# Potato Potahto in the FAO-GAEZ Productivity Measures?

### v1

@@ Introduction
where $Y_{res}=M\cdot Y$, $X_{res}^{*}=M\cdot X^{*}$, $\epsilon_{res}=M\cdot\epsilon$, $M\coloneqq I-W\cdot\left(W^{T}\cdot W\right)^{-1}\cdot W^{T}$ is the residual-maker matrix associated with the covariates, $I$ is the $J+1$ identity matrix and $W=\left[1,W_{1},\ldots,W_{J}\right]^{T}$ is the vector of covariates including the constant term.
@@
To partially identify $\beta$, we impose six assumptions. We also contextualize them to our empirical setting.
@@
$\beta\coloneqq\tilde{\beta}\cdot Var\left(\tilde{X}_{res}^{*}\right)\text{ and}X_{res}^{*}\coloneqq\dfrac{\tilde{X}_{res}^{*}}{Var\left(\tilde{X}_{res}^{*}\right)},$
@@
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Moreover, if the researcher is unwilling to impose Assumption 3, Appendix D offers an alternative partial identification strategy that does not require prior knowledge of the direction of the true effect. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect.
@@
Assumption 4 imposes that the measurement error terms in our proxies of the true variable of interest are uncorrelated with the unobservable economic shock after controlling for observable covariates. This assumption weakens the implicit identifying assumption in the empirical literature about agricultural productivity (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) since applied researchers impose that $U_{1}=U_{2}=0$. In our empirical application, it intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable.
@@
We now state our main result: partial identification of $\beta$ (Equation (2)) using two proxies (Equation (3)) with nonclassical measurement error.999 As discussed in Section 2.2, the bounds in Proposition 1 account for model uncertainty in the sense that the researcher is not sure about the magnitude of the correlation between the true unobservable variable of interest and the proxies’ measurement error terms. Due to the uncertainty behind the underlying model, the researcher may prefer to use the bounds in Proposition 1 or the ones in Appendices C and D, which use different set of assumptions. Importantly, these bounds treat the populational distribution of the observable variables as known objects. Consequently, they do not account for sample uncertainty. This type of uncertainty will be addressed in Section 4.
@@
Second, the bounds in Equation (6) explore all the restrictions imposed by $\mathbb{E}\left[Y_{res}\right]$, $\mathbb{E}\left[Z_{1,res}\right]$, $\mathbb{E}\left[Z_{1,res}\right]$, $Var\left(Y_{res}\right)$, $Var\left(Z_{1,res}\right)$, $Var\left(Z_{2,res}\right)$, $Cov\left(Y_{res},Z_{1,res}\right)$, $Cov\left(Y_{res},Z_{2,res}\right)$, $Cov\left(Z_{1,res},Z_{2,res}\right)$ on the data generating process. Since they do not use restrictions imposed by all the moments of the data distribution, they are not sharp as defined by Canay and Shaikh (2017). However, in a linear model context, most researchers are willing to impose restrictions, at most, on the second moments of the underlying data-generating process, and our identification strategy exhausts the information contained in these types of moments. [^10] A similar approach was adopted by Mogstad et al. (2018) in a different setting. Focusing on an IV problem, their partial identification results (Mogstad et al., 2018, Proposition 3) explore all the information contained in the conditional expectation of the outcome variable given the instrument, but do not explore higher-order moments.
@@
where $y_{jt}$ is the employment share of the manufacturing sector in municipality $j$ and year $t$, and $\Delta A_{jt}^{soy}$ is the soybean productivity shock. Bustos et al. (2016) explores an arguably exogenous shock in soybean productivity: the introduction of genetically modified seeds in 2003, i.e., they use this increase in productivity as an exogenous change in soybean productivity.
@@
We begin by discussing the differences in soybean productivity in versions 3 and 4. We are able to reproduce the numbers of productivity using Bustos et al. (2016)’s replication package. The correlation between the change in productivity variable in the replication package and the one generated by us is of 0.99.121212There are many reasons for differences in the computed productivity. In the replication package, there is no mention of how the variables were transformed from raster and vector formats to the tabular format the authors use. We use the geometry of the minimum comparable areas that is present in the replication package in order to compute the average productivity within each municipality. The replication package does not have the raw soybean productivity data, but we know of no different vintages of the FAO-GAEZ version 3 and, therefore, we have no reason to believe that the data is different. The differences then are likely to occur from different ways of assigning pixels of soybean productivity to the municipality geometry. For example: are all touched pixels considered? If yes, how? What is the projection used to make compatible the raster file of FAO-GAEZ with the vector file of the municipalities? There are also small holes in some municipalities in the Northeast region that we do not know if it was treated somehow after the data were generated. To ensure that all the differences we point out are coming from different FAO-GAEZ versions and not different data handling, we compare the results across our constructed productivity variables.
@@ Appendix D No Prior Knowledge of the Target Parameter’s Sign
Moreover, these bounds explore all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$.

### v2

@@ Introduction
where $Y_{res}\coloneqq Y-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Y\right]$, $X_{res}^{*}\coloneqq X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$, $\epsilon_{res}\coloneqq\epsilon-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot\epsilon\right]$, and $W=\left[1,W_{1},\ldots,W_{J}\right]$ is the row vector of covariates including the constant term.
@@
To partially identify $\beta$, we impose six assumptions, contextualizing them to our empirical setting. For brevity, we impose assumptions directly on the residualized variables in Equations (2) and (3). In Appendix H, we impose assumptions on the primitives of our model (Equation (1)) that imply the assumptions described in this section.
@@
$\beta\coloneqq\tilde{\beta}\cdot\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}\text{ and }X_{res}^{*}\coloneqq\dfrac{\tilde{X}_{res}^{*}}{\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}},$
@@
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Moreover, if the researcher is unwilling to impose Assumption 3, Appendix D offers an alternative partial identification strategy that does not require prior knowledge of the direction of the true effect. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign.
@@
Assumption 4 imposes that the measurement error terms in our proxies of the true variable of interest are uncorrelated with the unobservable economic shock after controlling for observable covariates. [^8] This assumptions is implied by $Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*},Z_{1},Z_{2}\right\}$ as discussed in Appendix H. This assumption weakens the implicit identifying assumption in the empirical literature about agricultural productivity (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) since applied researchers impose that $U_{1}=U_{2}=0$. In our empirical application, it intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable.
@@
The measurement error terms are positively correlated with each other, i.e., $Cov\left(U_{1},U_{2}\right)\geq 0$. We name this assumption nonclassical measurement error because the covariances between the true variable of interest and the measurement error terms are unrestricted, i.e., $Cov\left(X_{res}^{*},U_{k}\right)\in\mathbb{R}$ for any $k\in\left\{1,2\right\}$.
@@
We now state our main result: partial identification of $\beta$ (Equation (2)) using two proxies (Equation (3)) with nonclassical measurement error.101010 As discussed in Section 2.2, the bounds in Proposition 1 account for model uncertainty in the sense that the researcher is not sure about the magnitude of the correlation between the true unobservable variable of interest and the proxies’ measurement error terms. Due to the uncertainty behind the underlying model, the researcher may prefer to use the bounds in Proposition 1 or the ones in Appendices C and D, which use different set of assumptions. Importantly, these bounds treat the populational distribution of the observable variables as known objects. Consequently, they do not account for sample uncertainty. This type of uncertainty will be addressed in Section 4.
@@
Second, the bounds in Equation (6) exploit all the restrictions imposed by $\mathbb{E}\left[Y_{res}\right]$, $\mathbb{E}\left[Z_{1,res}\right]$, $\mathbb{E}\left[Z_{1,res}\right]$, $Var\left(Y_{res}\right)$, $Var\left(Z_{1,res}\right)$, $Var\left(Z_{2,res}\right)$, $Cov\left(Y_{res},Z_{1,res}\right)$, $Cov\left(Y_{res},Z_{2,res}\right)$, $Cov\left(Z_{1,res},Z_{2,res}\right)$ on the data generating process. Since they do not use restrictions imposed by all the moments of the data distribution, they are not sharp as defined by Canay and Shaikh (2017). However, in a linear model context, most researchers are willing to impose restrictions, at most, on the second moments of the underlying data-generating process, and our identification strategy exhausts the information contained in these types of moments. [^11] A similar approach was adopted by Mogstad et al. (2018) in a different setting. Focusing on an IV problem, their partial identification results (Mogstad et al., 2018, Proposition 3) exploit all the information contained in the conditional expectation of the outcome variable given the instrument, but do not exploit higher-order moments.
@@
where $y_{jt}$ is the employment share of the manufacturing sector in municipality $j$ and year $t$, and $\Delta A_{jt}^{soy}$ is the soybean productivity shock. Bustos et al. (2016) exploits an arguably exogenous shock in soybean productivity: the introduction of genetically modified seeds in 2003, i.e., they use this increase in productivity as an exogenous change in soybean productivity.
@@
We begin by discussing the differences in soybean productivity in versions 3 and 4. We are able to reproduce the numbers of productivity using Bustos et al. (2016)’s replication package. The correlation between the change in productivity variable in the replication package and the one generated by us is of 0.99.131313There are many reasons for differences in the computed productivity. In the replication package, there is no mention of how the variables were transformed from raster and vector formats to the tabular format the authors use. We use the geometry of the minimum comparable areas that is present in the replication package in order to compute the average productivity within each municipality. The replication package does not have the raw soybean productivity data, but we know of no different vintages of the FAO-GAEZ version 3 and, therefore, we have no reason to believe that the data is different. The differences then are likely to occur from different ways of assigning pixels of soybean productivity to the municipality geometry. For example: are all touched pixels considered? If yes, how? What is the projection used to make compatible the raster file of FAO-GAEZ with the vector file of the municipalities? There are also small holes in some municipalities in the Northeast region that we do not know if it was treated somehow after the data were generated. To ensure that all the differences we point out are coming from different FAO-GAEZ versions and not different data handling, we compare the results across our constructed productivity variables.
@@ Appendix D No Prior Knowledge of the Target Parameter’s Sign
Moreover, these bounds exploit all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$.
@@ Appendix G A Discussion on Partial Identification
Appendix H Assumptions on Model Primitives
In this appendix, we describe five assumptions on the primitives of our model that imply Assumptions 1-5 in Section 3.1. The primitives of our model are the random variables $Y$, $W_{1}$, …, $W_{J}$, $X^{*}$, $\epsilon$, $Z_{1}$ and $Z_{2}$, and the following outcome equation:
$Y=\alpha_{0}+\sum_{j=1}^{J}\alpha_{j}\cdot W_{j}+\beta\cdot X^{*}+\epsilon,$ (H.1)
Using the random variables in our model primitives, we define five additional random variables:
$X_{res}^{*}\coloneqq X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right],$ (H.2)
$Z_{k,res}\coloneqq Z_{k}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{k}\right]$ (H.3)
and
$U_{k}\coloneqq Z_{k,res}-X_{res}^{*}$ (H.4)
for any $k\in\left\{1,2\right\}$, where $W=\left[1,W_{1},\ldots,W_{J}\right]$ is the row vector of covariates including the constant term.
Our first assumption on the primitives of our model imposes that the economic shock is exogenous.
Assumption H.1
$Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*}\right\}$.
This assumption implies that
$\epsilon_{res}$ $\coloneqq\epsilon-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot\epsilon\right]$
$=\epsilon$
and that
$Cov\left(X_{res}^{*},\epsilon_{res}\right)$ $=Cov\left(X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right],\epsilon\right)$
$=Cov\left(X^{*},\epsilon\right)-Cov\left(W,\epsilon\right)\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
$=0.$
Hence, Assumption H.1 implies Assumption 1.
Our second assumption on the primitives of our model imposes a restriction on the variance of true unobservable measure.
Assumption H.2
$Var\left(X^{*}\right)=1+Var\left(W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right)$.
Note that
$Var\left(X^{*}\right)$ $=Var\left(X_{res}^{*}+W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right)$
$=Var\left(X_{res}^{*}\right)+Var\left(W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right)$
$\hskip 20.0pt+2\cdot Cov\left(X_{res}^{*},W\right)\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
$=Var\left(X_{res}^{*}\right)+Var\left(W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right).$
Hence, Assumption H.2 implies Assumption 2
Our third assumption on the primitives of our model is Assumption 3 directly.
Our fourth assumption on the primitives of our model imposes that the proxy variables are exogenous.
Assumption H.3
$Cov\left(Z_{k},\epsilon\right)=0$ for any $k\in\left\{1,2\right\}$.
Fix $k\in\left\{1,2\right\}$ arbitrarily and note that
$Cov\left(U_{k},\epsilon_{res}\right)$ $=Cov\left(U_{k},\epsilon\right)$
$=Cov\left(Z_{k,res}-X_{res}^{*},\epsilon\right)$
$=Cov\left(Z_{k}-W\cdot\delta_{k}-X^{*}+W\cdot\delta_{X},\epsilon\right)$
where $\hskip 20.0pt\text{where }\delta_{k}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{k}\right]$
and $\hskip 20.0pt\text{and }\delta_{X}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
$=Cov\left(Z_{k}-X^{*}-W\cdot\left(\delta_{k}-\delta_{X}\right),\epsilon\right)$
$=Cov\left(Z_{k},\epsilon\right)-Cov\left(X^{*},\epsilon\right)-Cov\left(W,\epsilon\right)\cdot\left(\delta_{k}-\delta_{X}\right)$
$=0.$
Hence, Assumptions H.1 and H.3 implies Assumption 4.
Lastly, since $U_{1}$ and $U_{2}$ are defined based on model primitives, our fifth assumption is Assumption 5 directly.